9.4. Let us again consider the elliptic curve y*x 2x 2 mod 17. Why are all points primitive elements?
Answer #1

Answer:

  • Consider the elliptic curve y^2 = x^3+2x + 2
  • An elliptic curve EK clear over a field k of point not equal to 2 or 3 is the set of explanation (X,Y) belong to k to the equation

Y^2 = x^3+ax+b, where a, b belongs to K

  • The points on elliptic curve^ (I.e. a point at infinity) structure a group below a certain addition law.
  • A primitive point P is only a generator of this collection; every elements of the group can be spoken as P+P+P+ . . . . . . . . . . . . . . . . . . . . . . . +P (k times) for a number of k.
  • if the elliptic curve has a prime number of points , then all its points ( except the point at infinity) are primitive, but in general, the elliptic can or may not have a primitive point...